# Security of Language Models for Code: A Systematic Literature Review

### v1

@@ 4.3.4. Summary of Existing Adversarial Attacks
Bielik et al. (Bielik and Vechev, 2020) 2020 ICML White-box attack LSTM DeepTyper GCN Type prediction
@@
CARROTA (Zhang et al., 2022a) 2022 TOSEM White-box attack GRU LSTM ASTNN LSCNN TBCNN CodeBERT CDLH Function classification Clone detection Vulnerability detection
@@
Anand et al. (Anand et al., 2021a) 2021 CoRR Black-box attack SketchAdapt SAPS Seq2Tree Code generation
ACCENT (Zhou et al., 2022) 2022 TOSEM Black-box attack LSTM GNN CSCG Rencos Code summarization
@@
Li et al. (Li et al., 2022a) 2022 ICSE Black-box attack DL-CAIS PbNN Authorship attribution
@@
FuncFooler (Jia et al., 2022) 2022 ArXiv Black-box attack RNN PV-DM Transformer Binary similarity detection
@@
RNNS (Zhang et al., 2023c) 2023 ArXiv Black-box attack CodeBERT, CodeT5 GraphCodeBERT Clone detection Vulnerability detection Authorship attribution Function classification
SHIELD (Abuhamad et al., 2023) 2023 CoRR Black-box attack CNN, CSFS, DL-CAIS Authorship attribution
@@
RADAR (Yang et al., 2024e) 2024 TOSEM Black-box attack CodeGPT PLBART, CodeT5 Code generation
AWAL et al. (Awal et al., 2024) 2024 CoRR Black-box attack Logistic Regression Multi-Layered Perceptron Gradient Boosting Classifier Decision Tree, AdaBoost Random Forest, Bagging Defect detection Clone detection Code classification
DeceptPrompt (Wu et al., 2023) 2023 CoRR Black-box attack CodeLlama, StarChat WizardCoder Code generation Vulnerability detection
STRUCK (Zhang et al., 2024d) 2024 CollaborateCom Black-box attack LSTM, GRU, GCN GGNN, CodeBERT GraphCodeBERT UniXcoder Function classification Defect detection
@@
CARL (Yao et al., 2024) 2024 TOSEM Black-box attack CodeBERT CodeT5 UniXcoder GraphCodeBERT CodeGeeX2 Code summarization Code translation Code refinement
@@ 5. Answer to RQ2: Defenses on CodeLMs
5.1. Research Directions in Defenses against CodeLMs.
As shown in the right part of Figure 4, we construct a taxonomy based on the 12 research papers collected on defenses against CodeLMs. Our taxonomy includes two categories of research directions: backdoor defenses and adversarial defenses. There are only 4 papers that focus on studying backdoor defense in CodeLMs and 8 papers that are dedicated to researching adversarial attacks in CodeLMs. Furthermore, based on different defense scenarios, backdoor defenses can be subdivided into pre-training defenses, in-training defenses and post-training defenses, while adversarial defenses can be categorized into adversarial training, model modification and additional models. In backdoor defenses, there are only 2 papers on pre-training defenses, 1 paper on in-training defenses, and 1 paper on post-training defenses. Compared to the research on backdoor attacks against CodeLMs, research on backdoor defenses is still in its early stages. In adversarial defenses, there are only 4 papers on adversarial training, 1 paper on model modification, and 3 papers on additional models. Similarly, adversarial defenses for CodeLMs are still in the early stages.
@@ 5.2.1. Overview of Backdoor Defenses against CodeLMs
In backdoor attacks on CodeLM, attackers can deploy various methods at different stages of the CodeLM learning process and trigger the implanted backdoors during the model’s application phase. As a result, researchers have designed defense methods tailored to the characteristics of backdoor attacks, aiming to be applicable at different stages of the learning process. These methods seek to eliminate or reduce the success rate of backdoor attacks, thereby enhancing the security of CodeLM. Based on different defense objectives, existing backdoor defense methods can be categorized into pre-training defenses, in-training defenses and post-training defenses. Pre-training defenses focus on detecting and removing poisoned samples before training. In-training defenses aim to prevent backdoor insertion during the model training phase. Post-training defenses are applied after model training is completed.
5.2.2. Pre-training Defenses
@@
Ramakrishnan et al. (Ramakrishnan and Albarghouthi, 2022) propose a backdoor detection defense method for CodeLMs based on the spectrum method (Tran et al., 2018). They argue that the spectrum method, which uses only the right singular vectors for data sorting, does not always satisfy spectrum separability. Therefore, they replace it with the previous right singular vectors. Additionally, they use the encoder output vectors or context vectors based on two common CodeLM architectures, Seq2Seq and Code2Seq, to detect toxic samples. Experiments demonstrate that the proposed backdoor defense method effectively detects low poisoning rates of toxic samples and can eliminate or reduce the attack success rate.
@@
Li et al. (Li et al., 2024b) introduce an effective backdoor detection defense method called CodeDetector, which detects toxic samples in training data without losing clean samples. CodeDetector focuses on detecting triggers for defense and assumes that attackers may hide triggers within important labels. It utilizes the integrated gradients technique (Sundararajan et al., 2017) to extract all important labels from the training data. Subsequently, it identifies abnormal labels that significantly impact the model’s performance and treats them as potential triggers. All data containing these triggers is then flagged as toxic samples. Experiments show that compared to baseline methods such as syntax checkers based on program analysis tools, ONION (Qi et al., 2021), clustering methods, and spectrum methods relying on outlier detection, CodeDetector achieves a 35.95% higher recall and a 16.1% lower false positive rate. These results indicate that CodeDetector can detect more toxic samples while losing fewer clean samples.
5.2.3. In-training Defenses
@@
Yang et al. (Yang et al., 2024d) propose a backdoor detection defense method for CodeLMs called Deceptive Cross-Entropy (DeCE) to enhance the security of CodeLMs against backdoor attacks. The method leverages the phenomenon of “early learning,” where models are initially less sensitive to backdoor triggers but gradually become more susceptible to them, leading to overfitting. DeCE tackles this issue by employing a blending process that combines the model’s predicted probability distribution with a deceptive distribution, which is gradually introduced during training. Additionally, label smoothing is applied to the original labels to prevent the model from becoming overly confident and reduce the risk of overfitting to backdoor triggers. Through extensive experimentation on various models and code synthesis tasks, DeCE demonstrates superior performance in defending against backdoor attacks while maintaining model accuracy on clean datasets, outperforming existing defense techniques such as BKI (Chen and Dai, 2021) and In-trust Loss.
5.2.4. Post-training Defenses
@@ 5.2.5. Summary of Backdoor Defenses
CodePoisoner (Li et al., 2024b) 2024 TOSEM Pre-trainig defense LSTM TextCNN Transformer CodeBERT Defect detection Clone detection Code repair
DeCE et al. (Yang et al., 2024d) 2024 ArXiv In-training defense CodeBERT GraphCodeBERT CodeGen CodeT5 CodeT5p Code generation Code repair
OSEQL (Hussain et al., 2023a) 2023 ArXiv Post-training defense CodeBERT CodeT5 PLBART RoBARTa BART Defect detection Clone detection
@@ 5.3.1. Overview of Adversarial Defenses against CodeLMs
For the aforementioned adversarial attack methods, researchers have proposed various adversarial defense methods to enhance the model’s robustness or eliminate the impact of adversarial samples on the model. There are primarily three methods in adversarial defense: (1) Input modification: modifying the model training process or altering input data samples. Adversarial examples are usually introduced to the model for adversarial training to enhance its robustness. (2) Model modification: adding more sub-networks, modifying activation functions, or changing loss functions within the model. (3) Additional models: using external models as additional networks when dealing with unseen data samples for classification. The first method focuses on data samples input to the model, while the other two methods are more concerned with the model itself. The latter two methods can be further categorized into model defense and data detection. The goal of model defense is to enable the model to correctly identify adversarial samples, while data detection involves detecting adversarial samples and issuing warnings, leading to the model rejecting further processing of adversarial inputs. Currently, adversarial defense methods for CodeLMs mainly concentrate on the first method. Next, we will provide a detailed overview of relevant research on adversarial defense for CodeLMs.
5.3.2. Adversarial Training
@@
Li et al. (Li et al., 2022a) introduce an innovative framework called RoPGen, which strengthens the robustness of authorship attribution models. RoPGen combines data augmentation and gradient augmentation during adversarial training, making it challenging for attackers to manipulate and mimic the model. RoPGen employs automatic encoding style imitation attacks and automatic encoding style obfuscation attacks for data augmentation, effectively increasing the diversity of training data. RoPGen introduces perturbations to the gradients of deep neural networks for data augmentation, learning a robust CodeLM with diverse representations. RoPGen uses four datasets based on programming languages, such as C, C++, and Java, for evaluation. Experimental results indicate that RoPGen significantly improves the robustness of code authorship attribution models, reducing target attack success rates by an average of 22.8% and non-target attack success rates by 41.0%.
@@
Ramakrishnan et al. (Henkel et al., 2022) propose an adversarial training method for CodeLMs based on Madry et al.’s robust optimization objective (Madry et al., 2018). To reduce the complexity of adversarial training, the authors pre-generate program sketches by applying transformations to some programs. During training, only the generated program sketches are considered, and a gradient-based attack is used to search for the optimal parameters for the transformations. Experiments show that adversarial training with just one transformation of CodeLMs efficiently enhances model robustness.
@@
Gao et al. (Gao et al., 2023) propose the EverI defense method against discrete adversarial attacks on CodeLMs. EverI trains the model on the strongest adversarial examples, making it theoretically applicable to defend against all adversarial attack methods. For discrete adversarial attacks, EverI first constructs a transitive closure based on semantic-preserving transformations on the original input, identifying adversarial examples in all programs within the transitive closure that cause the maximum loss for the model. For other classical continuous adversarial attacks like DAMP (Yefet et al., 2020), EverI identifies the strongest adversarial examples by adding adversarial steps and random initialization of adversarial perturbations. For other discrete attacks like Imitator (Quiring et al., 2019), EverI increases the width and depth of the program transformation search tree to find the strongest adversarial examples. Experiments show that, compared to adversarial training (Madry et al., 2018) and outlier detection defense (Yefet et al., 2020), EverI is more effective in defending against discrete adversarial attacks.
@@ 5.3.3. Model Modification
He et al. (He and Vechev, 2023) propose a novel learning-based secure control method SVEN for CodeLMs in code generation tasks and test the model’s security using crafted adversarial examples. SVEN utilizes continuous vectors with specific attributes to guide CodeLMs toward the direction of the given attributes in program generation without modifying the model’s weights. SVEN’s training process enforces specialized loss terms on different regions of the code and optimizes these continuous vectors using a carefully crafted high-quality dataset. Experiments show that SVEN can achieve effective secure control, increasing the proportion of secure code generation to 92.3% for CodeGen models without compromising the functional correctness of the generated code.
@@ 5.3.4. Additional Models
Zhang et al. (Zhang et al., 2023e) first generate adversarial examples using white-box attacks on smaller code models and then explore their effectiveness for transfer adversarial attacks on code LLMs. Meanwhile, they propose leveraging LLMs’ in-context learning and the self-defense capability of understanding human instructions, as well as meta-prompting techniques (i.e., using the LLM to generate its own defense prompts) to achieve adversarial defense for large models. Experimental results show that these defense methods effectively reduce the attack success rate on GPT-3.5, GPT-4, Claude-Instant-1, Claude-2, and CodeLlama.
Yang et al. (Yang et al., 2024e) propose RADAR-Defense, a defense method to improve the robustness of CodeLMs by synthesizing high-quality method names from functional descriptions. Specifically, RADAR-Defense avoids altering the original model and instead focuses on improving input quality by generating semantically meaningful method names that are supplied to CodeLMs along with functional descriptions and other signature information. This retrieval-enhanced prompt training method effectively reinstates the performance of CodeLMs under adversarial attacks. Experiments demonstrate that RADAR-Defense is able to restore the CodeBLEU and Pass@1 scores in various CodeLMs, such as CodeGPT, PLBART, CodeT5, Replit, CodeGen, and CodeT5+, thereby significantly improving model robustness without requiring retraining.

### v2

@@ 4.3.4. Summary of Adversarial Attacks
Bielik et al. (Bielik and Vechev, 2020) 2020 ICML White-box attack LSTM GCN DeepTyper Type prediction
@@
Anand et al. (Anand et al., 2021a) 2021 ArXiv Black-box attack SketchAdapt SAPS, Seq2Tree Code generation
ACCENT (Zhou et al., 2022) 2022 TOSEM Black-box attack LSTM, GNN CSCG, Rencos Code summarization
@@
CARROTA (Zhang et al., 2022a) 2022 TOSEM White-box attack GRU, LSTM ASTNN, LSCNN TBCNN, CDLH CodeBERT Function classification Clone detection Vulnerability detection
Li et al. (Li et al., 2022a) 2022 ICSE Black-box attack DL-CAIS, PbNN Authorship attribution
@@
FuncFooler (Jia et al., 2022) 2022 ArXiv Black-box attack RNN, PV-DM Transformer Binary similarity detection
@@
RNNS (Zhang et al., 2023a) 2023 ArXiv Black-box attack CodeBERT, CodeT5 GraphCodeBERT Clone detection Vulnerability detection Authorship attribution Function classification
SHIELD (Abuhamad et al., 2023) 2023 ArXiv Black-box attack CNN, CSFS, DL-CAIS Authorship attribution
@@
RADAR-Attack (Yang et al., 2024g) 2024 TOSEM Black-box attack CodeGPT PLBART, CodeT5 Code generation
Awal et al. (Awal et al., 2024) 2024 ArXiv Black-box attack Logistic Regression Multi-Layered Perceptron Gradient Boosting Classifier Decision Tree, AdaBoost Random Forest, Bagging Defect detection Clone detection Code classification
DeceptPrompt (Wu et al., 2023) 2023 ArXiv Black-box attack CodeLlama, StarChat WizardCoder Code generation Vulnerability detection
STRUCK (Zhang et al., 2023d) 2024 CollaborateCom Black-box attack LSTM, GRU, GCN GGNN, CodeBERT GraphCodeBERT UniXcoder Function classification Defect detection
@@
CARL (Yao et al., 2025) 2024 TOSEM Black-box attack CodeBERT, CodeT5 UniXcoder GraphCodeBERT CodeGeeX2 Code summarization Code translation Code refinement
@@ 5. Answer to RQ2: Defenses on CodeLMs
5.1. Research Directions in Defenses against CodeLMs
As shown in the right part of Figure 4, we construct a taxonomy based on the 13 research papers collected on defenses against CodeLMs. Our taxonomy includes two categories of research directions: backdoor defenses and adversarial defenses. There are only 4 papers that focus on studying backdoor defense in CodeLMs and 9 papers that are dedicated to researching adversarial attacks in CodeLMs. Furthermore, based on different defense scenarios, backdoor defenses can be subdivided into pre-training defenses, in-training defenses and post-training defenses, while adversarial defenses can be categorized into adversarial training, model modification and additional models. In backdoor defenses, there are only 2 papers on pre-training defenses, 1 paper on in-training defenses, and 1 paper on post-training defenses. Compared to the research on backdoor attacks against CodeLMs, research on backdoor defenses is still in its early stages. In adversarial defenses, there are only 5 papers on input modification, 1 paper on model modification, and 3 papers on additional models. Similarly, adversarial defenses for CodeLMs are still in the early stages.
@@ 5.2.1. Overview of Backdoor Defenses against CodeLMs
In backdoor attacks on CodeLM, attackers can deploy various methods at different stages of the CodeLM learning process and trigger the implanted backdoors during the model’s application phase. As a result, researchers have designed defense methods tailored to the characteristics of backdoor attacks, aiming to be applicable at different stages of the learning process. These methods seek to eliminate or reduce the success rate of backdoor attacks, thereby enhancing the security of CodeLM. Based on different defense objectives, existing backdoor defense methods can be categorized into pre-training defenses, in-training defenses and post-training defenses. Pre-training defenses focus on detecting and removing poisoned samples before training. They typically employ data filtering or anomaly detection techniques to identify potential backdoor samples in the dataset. In-training defenses aim to prevent backdoor insertion during the model training phase. They usually modify the training process to make the model less sensitive to the influence of poisoned samples during training. Post-training defenses are applied after model training is completed. They often utilize techniques such as model unlearning or input filtering to ensure model security.
5.2.2. Pre-training Backdoor Defenses
@@
Ramakrishnan et al. (Ramakrishnan and Albarghouthi, 2022) propose a backdoor detection defense method for CodeLMs based on the spectrum method (Tran et al., 2018). They argue that the spectrum method, which uses only the right singular vectors for data sorting, does not always satisfy $\varepsilon$-spectrum separability. In other words, the clean samples distribution and the poisoned samples distribution cannot be effectively separated in spectral space under the error tolerance controlled by $\varepsilon$. Therefore, they replace it with the top k right singular vectors. Additionally, they use the encoder output vectors or context vectors based on two common CodeLM architectures, Seq2Seq and Code2Seq, to detect poisoned samples. Experiments demonstrate that the proposed backdoor defense method effectively detects low poisoning rates of poisoned samples and can eliminate or reduce the attack success rate.
@@
Li et al. (Li et al., 2024b) introduce a backdoor detection defense method called CodeDetector, which detects poisoned samples in training data without losing clean samples. CodeDetector focuses on detecting triggers for defense and assumes that attackers may hide triggers within important labels. It utilizes the integrated gradients technique (Sundararajan et al., 2017) to extract all important labels from the training data. Subsequently, it identifies abnormal labels that significantly impact the model’s performance and treats them as potential triggers. All data containing these triggers is then flagged as poisoned samples. Experiments show that compared to baseline methods such as syntax checkers based on program analysis tools, ONION (Qi et al., 2021), clustering methods, and spectrum methods relying on outlier detection, CodeDetector achieves a 35.95% higher recall and a 16.1% lower false positive rate.
5.2.3. In-training Backdoor Defenses
@@
Yang et al. (Yang et al., 2024f) propose a backdoor detection defense method for CodeLMs called Deceptive Cross-Entropy (DeCE) to enhance the security of CodeLMs against backdoor attacks. The method leverages the phenomenon of “early learning”, where models are initially less sensitive to backdoor triggers but gradually become more susceptible to them, leading to overfitting. DeCE tackles this issue by employing a blending process that combines the model’s predicted probability distribution with a deceptive distribution, which is gradually introduced during training. Additionally, label smoothing is applied to the original labels to prevent the model from becoming overly confident and reduce the risk of overfitting to backdoor triggers. Through extensive experimentation on various models and code synthesis tasks, DeCE demonstrates superior performance in defending against backdoor attacks while maintaining model accuracy on clean datasets.
5.2.4. Post-training Backdoor Defenses
@@ 5.2.5. Summary of Backdoor Defenses
CodeDetector (Li et al., 2024b) 2024 TOSEM Pre-trainig defense LSTM, TextCNN Transformer CodeBERT Defect detection Clone detection Code repair
DeCE (Yang et al., 2024f) 2024 ArXiv In-training defense CodeBERT, CodeGen CodeT5, CodeT5p GraphCodeBERT Code generation Code repair
OSEQL (Hussain et al., 2023a) 2023 ArXiv Post-training defense CodeBERT, CodeT5 BART, PLBART RoBARTa Defect detection Clone detection
@@ 5.3.1. Overview of Adversarial Defenses against CodeLMs
For the aforementioned adversarial attack methods, researchers have proposed various adversarial defense methods to enhance the model’s robustness or eliminate the impact of adversarial samples on the model. There are primarily three methods in adversarial defense: (1) Input modification: modifying the model training process, altering input data samples or optimizing the training process. The core idea is to proactively detect and mitigate the impact of adversarial samples or strengthen the model’s resistance to abnormal inputs through adversarial training. Typically, adversarial examples are incorporated into the training set, allowing the model to adapt to adversarial perturbations during the learning process and improving its stability against attacks. (2) Model modification: adding more sub-networks, modifying activation functions, changing loss functions within the model, or introducing regularization techniques. Robustness-optimized loss functions are commonly employed to improve classification accuracy in adversarial regions. Additionally, adjusting activation functions can increase the model’s sensitivity to adversarial samples, thereby reducing attack success rates. (3) Additional models: using external models as additional networks when dealing with unseen data samples for classification. The core idea is to use additional detection mechanisms or adversarially trained networks to identify and process potential adversarial samples. Typically, an adversarial sample detector is placed before the original model to filter input data, ensuring that only normal inputs are processed by the model. The first method focuses on data samples input to the model, while the other two methods are more concerned with the model itself. The latter two methods can be further categorized into model defense and data detection. The goal of model defense is to enable the model to correctly identify adversarial samples, while data detection involves detecting adversarial samples and issuing warnings, leading to the model rejecting further processing of adversarial inputs. Currently, adversarial defense methods for CodeLMs mainly concentrate on the first method. Next, we will provide a detailed overview of relevant research on adversarial defense for CodeLMs.
5.3.2. Input Modification
@@
Li et al. (Li et al., 2022a) introduce a framework called RoPGen, which strengthens the robustness of authorship attribution models. RoPGen combines data augmentation and gradient augmentation during adversarial training, making it challenging for attackers to manipulate and mimic the model. RoPGen employs automatic encoding style imitation attacks and automatic encoding style obfuscation attacks for data augmentation, effectively increasing the diversity of training data. RoPGen introduces perturbations to the gradients of deep neural networks for data augmentation, learning a robust CodeLM with diverse representations. RoPGen uses four datasets based on programming languages, such as C, C++, and Java, for evaluation. Experimental results indicate that RoPGen significantly improves the robustness of code authorship attribution models, reducing target attack success rates by an average of 22.8% and non-target attack success rates by 41.0%.
@@
To reduce the complexity of adversarial training, Henkel et al. (Henkel et al., 2022) propose an adversarial training method for CodeLMs based on Madry et al.’s robust optimization objective (Madry et al., 2018). They pre-generate program sketches by applying transformations to some programs. During training, only the generated program sketches are considered, and a gradient-based attack is used to search for the optimal parameters for the transformations. Experiments show that adversarial training with just one transformation of CodeLMs efficiently enhances model robustness.
To enhance the robustness of source code processing models, Zhang et al. (Zhang et al., 2022a) further proposed the adversarial training method CARROTT. CARROTT utilizes adversarial examples generated by the attack method CARROTA to periodically augment the training set and introduce adversarial perturbations during the training process, thus systematically enhancing the model’s robustness to adversarial attacks. Experimental evaluations show that models trained with CARROTT exhibit an average robustness improvement of 5.3 times across multiple tasks (such as code functionality classification, code clone detection, and defect prediction), significantly outperforming MHM, which achieves only a 1.7× improvement.
@@
Furthermore, Gao et al. (Gao et al., 2023) propose the EverI defense method against discrete adversarial attacks on CodeLMs. EverI trains the model on the strongest adversarial examples, making it theoretically applicable to defend against all adversarial attack methods. For discrete adversarial attacks, EverI first constructs a transitive closure based on semantic-preserving transformations on the original input, identifying adversarial examples in all programs within the transitive closure that cause the maximum loss for the model. For other classical continuous adversarial attacks like DAMP (Yefet et al., 2020), EverI identifies the strongest adversarial examples by adding adversarial steps and random initialization of adversarial perturbations. For other discrete attacks like Imitator (Quiring et al., 2019), EverI increases the width and depth of the program transformation search tree to find the strongest adversarial examples. Experiments show that, compared to adversarial training (Madry et al., 2018) and outlier detection defense (Yefet et al., 2020), EverI is more effective in defending against discrete adversarial attacks.
@@ 5.3.3. Model Modification
He et al. (He and Vechev, 2023) propose a learning-based secure control method SVEN for CodeLMs in code generation tasks and test the model’s security using crafted adversarial examples. SVEN utilizes continuous vectors with specific attributes to guide CodeLMs toward the direction of the given attributes in program generation without modifying the model’s weights. SVEN’s training process enforces specialized loss terms on different regions of the code and optimizes these continuous vectors using a carefully crafted high-quality dataset. Experiments show that SVEN can achieve effective secure control, increasing the proportion of secure code generation to 92.3% for CodeGen models without compromising the functional correctness of the generated code.
@@ 5.3.4. Additional Models
Zhang et al. (Zhang et al., 2023c) first generate adversarial examples using white-box attacks on smaller code models and then explore their effectiveness for transfer adversarial attacks on code LLMs. Meanwhile, they propose leveraging LLMs’ in-context learning and the self-defense capability of understanding human instructions, as well as meta-prompting techniques (i.e., using the LLM to generate its own defense prompts) to achieve adversarial defense for large models. Experimental results show that these defense methods effectively reduce the attack success rate on GPT-3.5, GPT-4, Claude-Instant-1, Claude-2, and CodeLlama.
Yang et al. (Yang et al., 2024g) propose RADAR-Defense, a defense method to improve the robustness of CodeLMs by synthesizing high-quality method names from functional descriptions. Specifically, RADAR-Defense avoids altering the original model and instead focuses on improving input quality by generating semantically meaningful method names that are supplied to CodeLMs along with functional descriptions and other signature information. This retrieval-enhanced prompt training method effectively reinstates the performance of CodeLMs under adversarial attacks. Experiments demonstrate that RADAR-Defense is able to restore the CodeBLEU and Pass@1 scores in various CodeLMs, such as CodeGPT, PLBART, CodeT5, Replit, CodeGen, and CodeT5+, thereby significantly improving model robustness without requiring retraining.